solve each of the following initial value problems

  1. (2x-5y)dx+(4x-y)dy=0 y(1)=4
  2. (3x2+9xy+5y2)dx-(6x2+4xy)dy=0 y(1)=4
1
Expert's answer
2021-07-13T08:50:41-0400

1.


"(2x-5y)dx+(4x-y)dy=0"

"2-5\dfrac{y}{x}+(4-\dfrac{y}{x})\dfrac{dy}{dx}=0"

Let "y=xv." Then "\dfrac{dy}{dx}=v+x\dfrac{dv}{dx}"


"2-5v+(4-v)(v+x\dfrac{dv}{dx})=0"

"2-5v+4v-v^2+x(4-v)\dfrac{dv}{dx}=0"

"\int\dfrac{v-4}{v^2+v-2}dv=-\int\dfrac{dx}{x}"

"\dfrac{v-4}{v^2+v-2}=\dfrac{A}{v-1}+\dfrac{B}{v+2}=\dfrac{A(v+2)+B(v-1)}{(v-1)(v+2)}"


"A(v+2)+B(v-1)=v-4"

"v=1: 3A=-3=>A=-1"

"v=-2: -3B=-6=>B=2"


"-\int\dfrac{dv}{v-1}+2\int\dfrac{dv}{v+2}=-\int\dfrac{dx}{x}"

"-\ln|v-1|+2\ln|v+2|=-\ln|x|+\ln C"

"\dfrac{x(v+2)^2}{(v-1)}=C"

"\dfrac{(y+2x)^2}{(y-x)}=C"

"y(1)=4"


"C=\dfrac{(4+2(1))^2}{(4-1)}=12"

"\dfrac{(y+2x)^2}{(y-x)}=12"

2.


"(3x^2+9xy+5y^2)dx-(6x^2+4xy)dy=0"

"3+9\dfrac{y}{x}+5(\dfrac{y}{x})^2-(6+4\dfrac{y}{x})\dfrac{dy}{dx}=0"

Let "y=xv." Then "\dfrac{dy}{dx}=v+x\dfrac{dv}{dx}"


"3+9v+5v^2-(6+4v)(v+x\dfrac{dv}{dx})=0"

"3+9v+5v^2-6v-4v^2-x(6+4v)\dfrac{dv}{dx}=0"

"\int\dfrac{6+4v}{v^2+3v+3}dv=\int\dfrac{dx}{x}"



"d(v^2+3v+3)=(2v+3)dv"


"2\ln(v^2+3v+3)=\ln|x|+\ln C"


"\dfrac{(v^2+3v+3)^2}{x}=C"

"\dfrac{(y^2+3xy+3x^2)^2}{x^5}=C"

"y(1)=4"


"\dfrac{(4^2+3(1)(4)+3(1)^2)^2}{1^5}=C"

"C=961"

"\dfrac{(y^2+3xy+3x^2)^2}{x^5}=961"